OpenStudy (anonymous):

Find the linear approximation of the function f(x)=^3root(1+x) at a=0 and use it to approximate ^3root(0.95) and ^3root(1.01).

OpenStudy (anonymous):

\[f(x)=\sqrt[3]{1+x}\] \[\sqrt[3]{0.95} \] \[\sqrt[3]{1.01}\]

OpenStudy (anonymous):

\[\sqrt[3]{1+x}\] has derivative \[\frac{1}{3\sqrt[3]{x+1}^2}\] at \(x=0\) you get the slope of \(\frac{1}{3}\) so the equation for your line is \[y=\frac{1}{3}x+1\]

OpenStudy (anonymous):

that is the "linearization" at \(x=0\) for \(1.01\) use \(x=.01\) and for \(.95\) use \(x=-.05\)